Tech Staff Meeting Notes November 14, 2005

Current Top Priorities are:Station Maintenance, Broadband Upgrade

There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about

John Filipcic will be leaving us as of November 18. The first half of the meeting was devoted to redistributing responsibilities. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything. The new list is here.

An abbreviated monthly meeting consumed the other 30 minutes.

Greg's October ToDo List


BB network upgrade.
Received the Intuicom radios from UUSS for testing. Steve is running the pdf's.

James provided this update. Got permission for a Radio link at Steele, MO. Need to return the borrowed receiver at Reelfoot and unfortunately only have an old one to replace it. Bob is submitting an NSF proposal to more than double the size of the network. We do siting, UNAVCO installs and owns the hardware. Allows 3 months/year staff time.

John Parker found a perfect site for a node at Mt Mitchell. Power, tower, AC, rack space, generator, DSL, etc. We'll be moving LRNC to Mt Mitchell. John will contact the prof at LRNC to see what they would like us to leave behind for display purposes. Need to look at what stations we don't really need (e.g. PLVA, PKNC).

Sensor swapping continues; Filipcic is taking the lead and keeping records.
CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. Filipcic needs to install. James brought back some nice pictures from CVTN to help figure out how to do it (e.g. the conduit is way too small).
Parker will be going to Charleston friday to figure out what's wrong with C2SC. Done.

Routine Maintenance. Jim knows what needs work.

I-40 Bridge.
W28 borehole sensor has a bad component.
A large piece of concrete fell and crushed the conduit to one of the sensors. Will need to pull new wires and install new flex conduit.
DAS #1 is at 50 sps and still requires frequent phone resets. Think it goes zombie when the link is dropped. It's the only one transmitting from the north side which probably makes the connection poor.
Need to get a summary report and some data to Sharam. Need to establish who wants what when.

Second Tier Priorities

FDL. We'll start keeping fdl notes in a separate fdl section. On hold til John F. gets back.

As Time Permits


Mitch will be in New Harmony for the "ribbon cutting" on November 3. Done.

MPHB. Withers needs to determine current orientation so we can pull it and orient correctly. Helicorders look strange. Need to verify the data.

Timing. Steiner has an easy way to generate a datastream that goes high when lock is lost. Need a mask to decide what streams to use. Need mechanism to produce alarms. BB stations more difficult. How do we automate the status stream or even get it back to Memphis for logging? (Save it to file with Scream?). Note, new version of firmware sends status to both flash and stream. Hence, information will be in gcf2ew logs, but how do we monitor that? Maybe not a huge deal with broadbands since both Bollwerk and Brewer regularly monitor. Would still be good to have a record though.

VBMS. Everything is in order and ready to go. Installation will be this fall depending on schedule of contractor. Greg may visit during installation if schedule permits.

Langston student projects continue. See Greg's list above.